OpenStudy (anonymous):

What exponential function represents the data in the table? x f(x) 2 25 3 125 4 625

OpenStudy (anonymous):

A) f(x) = x4 + 9 B) f(x) = 4x + 9 C) f(x) = x5 D) f(x) = 5x